Your marginal utility for consuming a can of Reeses smoked oysters is 28 utils and your marginal utility of consuming an order of French fries is 15 utils. A can of oysters cost $3, whereas, a bag of fries costs $1.50. 5. Are you maximizing your utility at the above consumption point? Why or why not? Why is the important part of answering this question. a. Given your answer in part a, how should you change consumption (if it should be changed) to maximize your utility? Why? Why is the important part of answering this question. b.
Answer #1

a) MRS= MUx/MUy= 28/15= 1.87

Px/Py= 3/1.5= 2

At utility maximisation , MRS= Px/Py

But this does not happen with the given combination, so you are not maximising utility.

b) MUx/MUy < Px/Py

To equate the two ratios, MUx has to increase and MUy has to decrease.

Decrease in consumption of oysters leads to increase MUx and increase in consumption of Fries decreases MUy.

Thus less of oysters and more of fries should be consumed.